OpenStudy (anonymous):

PLEASE HELP!!!!! (medal will be given to best answer) A mall employee is dressing a mannequin. There are 6 pairs of shoes, 4 types of jeans, and 8 sweaters. Using 1 of each, how many ways can the mannequin be dressed? Please explain how you'd solve this problem. Thanks! :)

Parth (parthkohli):

So, there are 6 ways to choose her shoes, 4 to choose the jeans and 8 for sweaters. We may multiply the numbers to get the number of ways.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow 6 \times 4 \times 8 = 6 \times 32 = 192 }\) 192 different ways.

OpenStudy (anonymous):

BTW - I pretty sure that's it's either 18C3 or 18P3 18C3 gives an answer of: 816 and 18P3 gives an answer of: 4896 Is this correct?

OpenStudy (anonymous):

Okay, I'm a bit confused - wouldn't multiplying not count the number of possible combination that you could have for every single combination?

OpenStudy (anonymous):

*combinations